OpenStudy (anonymous):

Determine whether each ordered pair is a solution of the given equation. 5y=2/3x+1; (6,1)

OpenStudy (anonymous):

the equation is, 5y=(2/3)x+1 putting (6,1) point, 5(1)=(2/3)6+1 =>5=2(2)+1 =>5=4+1 =>5=5 so the point satisfies the equation. so the point is a solution of the given equation.